The College of Business at Northeast College is accumulating data as a first step in the preparation of next year's budget development. One cost that is being looked at closely is administrative costs as a function of student credit hours. Data on administrative costs and credit hours for the past thirteen months are shown below:

Month Administrative Costs Credit Hours
July $ 129,551 500
August 82,813 245
September 225,730 1,522
October 216,494 1,130
November 258,313 1,439
December 184,699 1,242
January 219,337 1,469
February 245,150 1,503
March 209,562 1,194
April 191,975 1,253
May 250,228 1,490
June 170,618 550
July 128,317 445
Total $ 2,512,787 13,982
Average $ 193,291 1,076


The controller's office has analyzed the data and has given you the results from the regression analysis:

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.9326223
R Square 0.869784356
Adjusted R Square 0.857946570
Standard Error 19,994.71939
Observations 13
ANOVA
df SS MS F Significance F
Regression 1 29,374,584,992 29,374,584,992 -73.4752568 3.36855E-06
Residual 11 4,397,676,837 399,788,803.4
Total 12 33,772,261,829
Coefficients Standard Error t Stat P-value Lower 95% Upper 95% Lower 95.0% Upper 95.0%
Intercept 79,228.10053 14,416.13842 5.49579216 0.000187421 47,498.39380 110,957.8073 47,498.39380 110,957.8073
X Variable 1 106.0521880 12.37226111 8.571770925 3.36855E-06 78.82102495 133.283351 78.82102495 133.283351

  
Based on the results of the high-low analysis, the estimate of administrative costs in a month with 1,100 credit hours would be: (rounded to the nearest whole dollar)

  • $197,660.

  • $178,500.

  • $196,457.

  • $195,316.
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (225,730-82,813)/(1522-245) = 111.92

Fixed cost = 225,730 - (1522*111.92) = 55,388

Cost for 1100 hours = (1100*111.92) + 55388

= 178,500